Could someone please help solve this equation: p= sqrt (2p-1)
thanx

Answer by atif.muhammad(135)

p= sqrt (2p-1)
p^2 = 2p - 1
p^2 - 2p + 1 = 0
p^2 - p - p + 1 = 0
p(p-1) - 1(p-1) = 0
(p-1)(p-1) = 0
(p-1)^2 = 0
p-1 = 0
p=1
Your answer is p = 1.